OpenStudy (anonymous):

For which system of equations is (–2, 2) the solution? A. 3x – 5y = –16 –2x + 3y = 10 B. –4x + 2y = 6 2x – 3y = 10 C. 4x + 3y = 2 5x + 4y = –2 D. 6x + 5y = –2 –8x + 5y = –6

OpenStudy (anonymous):

substitute x=-2 into the equations

OpenStudy (anonymous):

if u get y=2,that will be ur correct equation

OpenStudy (anonymous):

Let's check one by one

OpenStudy (anonymous):

Option A...

OpenStudy (anonymous):

\[3x-5y=-16\]\[-2x+3y=10\]

OpenStudy (anonymous):



OpenStudy (anonymous):

substitute x=-2 into the equation

OpenStudy (anonymous):

\[3(-2)-5y=-16\]\[-2(-2)+3y=10\]

OpenStudy (anonymous):

if u get y=2,that equation is correct

OpenStudy (anonymous):

\[-6-5y=-16\]\[4+3y=10\]

OpenStudy (anonymous):

i know it not B bc it just doesn't have y=2

OpenStudy (anonymous):

\[-5y=-16+6\]\[3y=10-4\]\[-5y=-10\]\[3y=6\]\[y=2,y=2\]

OpenStudy (anonymous):

its C!

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

Option B...

OpenStudy (anonymous):

\[-4x+2y=6\]\[2x-3y=10\]\[-4(-2)+2y=6\]\[2(-2)-3y=10\]

OpenStudy (anonymous):

\[8+2y=6\]\[-4-3y=10\]

OpenStudy (anonymous):

its A?

OpenStudy (anonymous):

\[2y=6-8\]\[-3y=10+4\]

OpenStudy (anonymous):

yep :)

OpenStudy (anonymous):

thank you so much!

OpenStudy (anonymous):

yw!